# Integrability of a family of clean SYK models from the critical Ising chain

### v1

@@ Abstract
We establish the integrability of a family of SYK models with uniform $p$-body interactions. We derive the R-matrix and mutually commuting transfer matrices that generate the Hamiltonians of these models, and obtain their exact eigenspectra and eigenstates. Remarkably, the R-matrix is that of the critical transverse-field Ising chain. This work reveals an unexpected connection between the SYK model, central to many-body quantum chaos, and the critical Ising chain, a cornerstone of statistical mechanics.
@@ I Introduction
The Sachdev-Ye-Kitaev (SYK) model [1, 2, 3, 4, 5, 6] has emerged as a paradigmatic example of quantum many-body chaos, exhibiting maximal scrambling while remaining analytically tractable in the large-$N$ limit. The model consists of $N$ Majorana fermions with random all-to-all $p$-body interactions and saturates the chaos bound [7, 2, 8, 9], making it a valuable theoretical laboratory for studying quantum chaos and its connection to black hole physics [10, 11]. The essential role of disorder in the SYK model has motivated the search for simpler, disorder-free variants that retain the model’s key features [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. In particular, Witten showed that tensor models can reproduce the same large-$N$ limit as the SYK model without quenched disorder [15].
@@ II Clean SYK models
which immediately leads to the mutual commutativity of the supercharges:
@@ III Integrability from critical Ising chain
where we introduced two auxiliary Majorana fermions $\gamma_{a}$ and $\gamma_{b}$, and the same relation holds for the backward monodromy matrix $\overset{\leftarrow}{T}_{a}(u)$. Equation (14) can be proved using the Yang-Baxter equation (10) iteratively:
@@
Via the Jordan-Wigner transformation $\gamma_{2j-1}=\quantity(\prod_{l=1}^{j-1}Z_{l})X_{j}$ and $\gamma_{2j}=\quantity(\prod_{l=1}^{j-1}Z_{l})Y_{j}$, where $X_{j}$, $Y_{j}$, and $Z_{j}$ are the Pauli matrices acting on the $j$th site, we have
@@
Here, $P_{j,l}^{\pm}\equiv\gamma_{j}\pm\gamma_{l}$, and $\mathcal{U}^{\prime}\equiv\tau_{-}(-\mathrm{i})=\tau_{+}(-\mathrm{i})\gamma_{1}$. The twisted translation acts as
@@
where the indices are taken modulo $N$. Equations (26), (27) and (28) can be proved using a similar argument to that in Ref. [29]. We note that the action of the twisted translation (27) differs from that in Ref. [29]. This is because our monodromy matrix is constructed from the staggered choice of spectral parameters in Eqs. (12) and (13), whereas the monodromy matrix in Ref. [29] uses a uniform choice of spectral parameters.
@@
where $C_{\pm}=i(N\pm 1-1)/4$. The proof of Eq. (29) is given in Appendix E. Equation (29) means that the SYK charges with even/odd Majorana fermions commute with the critical Ising Hamiltonian $H_{\mathrm{Ising}}^{\mp}$, and then can be simultaneously diagonalized. Higher-order derivatives of the logarithm of the transfer matrix give the higher-order local conserved charges in the Ising chain [41, 42, 43, 44], which are all bilinear in Majorana fermions, as can be easily seen from Eqs. (31) and (32) below.
@@ IV Exact solution
where $\mathcal{I}^{+}\equiv\left\{k\in\frac{2\pi}{N}\quantity(\mathbb{Z}+\frac{1}{2}):0<k<\pi\right\}$ and $\mathcal{I}^{-}\equiv\left\{k\in\frac{2\pi}{N}\mathbb{Z}:0<k<\pi\right\}$. The fermionic creation operators are then defined by $f_{k}^{\dagger}$. These complex fermion operators satisfy the anti-commutation relations: $\{f_{k},f_{k^{\prime}}\}=\delta_{k,k^{\prime}}$, $\{f_{k},f_{k^{\prime}}\}=\{f^{\dagger}_{k},f^{\dagger}_{k^{\prime}}\}=0$. We also define the Majorana zero mode: $\chi_{0}\equiv\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\gamma_{j}$ and the Majorana $\pi$ mode: $\chi_{\pi}\equiv\frac{1}{\sqrt{N}}\sum_{j=1}^{N}(-1)^{j-1}\gamma_{j}$, satisfying $\chi_{0}^{2}=\chi_{\pi}^{2}=1$ and $\{\chi_{0},f_{k}\}=\{\chi_{\pi},f_{k}\}=\{\chi_{0},\chi_{\pi}\}=0$.
@@
Note that $\epsilon_{k}$ are the single-particle energies of the quadratic Hamiltonians: $H_{2}$ for $k\in\mathcal{I}^{+}$ and $\sqrt{N}\chi_{0}H_{3}$ for $k\in\mathcal{I}^{-}$ [13]. These quadratic forms enabled the diagonalization of $H_{N}^{(4)}$ and $(H_{N}^{(3)})^{2}$ with the complex fermions $f_{k}$ in [13]; here we extend this to all SYK charges $H_{N}^{(p)}$. The eigenvalues of $\tau_{\pm}(u)$ are thus given by $(\sqrt{N}s_{0})^{\delta_{\pm}}\prod_{k\in\mathcal{I}^{\pm}}(1-u\cot(k/2)s_{k})$ where $s_{k}\in\{+1,-1\}$, $\delta_{+}\equiv 0$ and $\delta_{-}\equiv 1$. The proof of Eqs. (31) and (32) is given in Appendix H. For odd $N$, $\chi_{\pi}$ commutes with $\tau_{+}(u)$. For even $N$, $\chi_{\pi}$ anticommutes with $\tau_{-}(u)$.
@@
The above exact solution means that we have also obtained the exact solution for hybrid models consisting of the SYK charges and the local charges in the critical Ising chain. Such models, with Hamiltonians that are linear combinations of long-range and short-range interaction terms, have also been investigated in the literature [56], where the SYK model and the Kitaev chain are coupled. Our results may provide new insights into the nature of such models.
@@ V Conclusion
We have established the complete integrability of the clean SYK models with $p$-body interactions by revealing their connection to the critical transverse field Ising chain. The key discovery is that the generalized SYK Hamiltonians emerge from transfer matrices constructed using the R-matrix of the critical Ising chain, which satisfies the Yang-Baxter equation. Our approach provides explicit solutions for all clean SYK models, extending the previously known results [13] for $H_{4}$ and $(H_{3})^{2}$ to the general $p$-body case $H_{p}$. The unexpected connection between the family of clean SYK models and the critical Ising chain places the former within the well-established framework of Yang-Baxter integrability.

### v2

@@ Abstract
We establish the integrability of a family of Sachdev-Ye-Kitaev (SYK) models with uniform $p$-body interactions. We derive the R-matrix and mutually commuting transfer matrices that generate the Hamiltonians of these models, and obtain their exact eigenspectra and eigenstates. Remarkably, the R-matrix is that of the critical transverse-field Ising chain. This work reveals an unexpected connection between the SYK model, central to many-body quantum chaos, and the critical Ising chain, a cornerstone of statistical mechanics.
@@ I Introduction
The Sachdev-Ye-Kitaev (SYK) model [1, 2, 3, 4, 5, 6, 7] has emerged as a paradigmatic example of quantum many-body chaos, exhibiting maximal scrambling while remaining analytically tractable in the large-$N$ limit. The model consists of $N$ Majorana fermions with random all-to-all $p$-body interactions and saturates the chaos bound [8, 2, 9, 10], making it a valuable theoretical laboratory for studying quantum chaos and its connection to black hole physics [11, 12]. Moreover, the model realizes strange metal behavior in the large-$N$ limit—a non-Fermi liquid phase relevant to cuprate superconductors [13, 6]. The essential role of disorder in the SYK model has motivated the search for simpler, disorder-free variants that retain the model’s key features [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. In particular, Witten showed that tensor models can reproduce the same large-$N$ limit as the SYK model without quenched disorder [17].
@@ II Clean SYK models
which immediately leads to the mutual commutativity of the SYK supercharges:
@@ III Integrability from critical Ising chain
where we introduced two auxiliary Majorana fermions $\gamma_{a}$ and $\gamma_{b}$, and the same relation holds for the backward monodromy matrix $\overset{\leftarrow}{T}_{a}(u)$. Equation (14) can be proved using the Yang-Baxter equation (10) repeatedly:
@@
The choice of the sign ($\pm$) corresponds to periodic (++) or anti-periodic (-) boundary conditions on the Majorana fermions. The Hamiltonian $H_{\mathrm{Ising}}^{+}$ commutes with the supercharges $H_{2p+1}$, whereas $H_{\mathrm{Ising}}^{-}$ commutes with the SYK Hamiltonians $H_{2p}$, which will be proved below. Via the Jordan-Wigner transformation $\gamma_{2j-1}=\quantity(\prod_{l=1}^{j-1}Z_{l})X_{j}$ and $\gamma_{2j}=\quantity(\prod_{l=1}^{j-1}Z_{l})Y_{j}$, where $X_{j}$, $Y_{j}$, and $Z_{j}$ are the Pauli matrices acting on the $j$th site, we have
@@
Here we defined $P_{j,l}^{\pm}\equiv\gamma_{j}\pm\gamma_{l}$. The twisted translation acts as
@@
where $\mathcal{U}^{\prime}\equiv\tau_{-}(-\mathrm{i})=\tau_{+}(-\mathrm{i})\gamma_{1}$ and the indices are taken modulo $N$. Equations (26), (27) and (28) can be proved using a similar argument to that in Ref. [31]. We note that the action of the twisted translation (27) differs from that in Ref. [31]. This is because our monodromy matrix is constructed from the staggered choice of spectral parameters in Eqs. (12) and (13), whereas the monodromy matrix in Ref. [31] uses a uniform choice of spectral parameters.
@@
where $C_{\pm}=\mathrm{i}(N\pm 1-1)/4$. The proof of Eq. (29) is given in Appendix E. Equation (29) means that the SYK charges with even/odd Majorana fermions commute with the critical Ising Hamiltonian $H_{\mathrm{Ising}}^{\mp}$, and then can be simultaneously diagonalized. Higher-order derivatives of the logarithm of the transfer matrix give the higher-order local conserved charges in the Ising chain [43, 44, 45, 46], which are all bilinear in Majorana fermions, as can be easily seen from Eqs. (31) and (32) below.
@@ IV Exact solution
where $\mathcal{I}^{+}\equiv\left\{k\in\frac{2\pi}{N}\quantity(\mathbb{Z}+\frac{1}{2}):0<k<\pi\right\}$ and $\mathcal{I}^{-}\equiv\left\{k\in\frac{2\pi}{N}\mathbb{Z}:0<k<\pi\right\}$. The fermionic creation operators are then defined by $f_{k}^{\dagger}$. These complex fermion operators satisfy the anti-commutation relations: $\{f_{k}^{\dagger},f_{k^{\prime}}\}=\delta_{k,k^{\prime}}$, $\{f_{k},f_{k^{\prime}}\}=\{f^{\dagger}_{k},f^{\dagger}_{k^{\prime}}\}=0$. We also define the Majorana zero mode: $\chi_{0}\equiv\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\gamma_{j}$ and the Majorana $\pi$ mode: $\chi_{\pi}\equiv\frac{1}{\sqrt{N}}\sum_{j=1}^{N}(-1)^{j-1}\gamma_{j}$, satisfying $\chi_{0}^{2}=\chi_{\pi}^{2}=1$ and $\{\chi_{0},\chi_{\pi}\}=0$. Also, they satisfy $\{\chi_{0},f_{k}\}=0$ for $k\in\mathcal{I}^{-}$ and $\{\chi_{\pi},f_{k}\}=0$ for $k\in\mathcal{I}^{(-)^{N+1}}$.
@@
Note that $\epsilon_{k}$ are the single-particle energies of the quadratic Hamiltonians: $H_{2}$ for $k\in\mathcal{I}^{+}$ and $\sqrt{N}\chi_{0}H_{3}$ for $k\in\mathcal{I}^{-}$ [15]. These quadratic forms enabled the diagonalization of $H_{N}^{(4)}$ and $(H_{N}^{(3)})^{2}$ with the complex fermions $f_{k}$ in [15]; here we extend this to all SYK charges $H_{N}^{(p)}$. The eigenvalues of $\tau_{\pm}(u)$ are thus given by $(\sqrt{N}s_{0})^{\delta_{\pm}}\prod_{k\in\mathcal{I}^{\pm}}(1-u\cot(k/2)s_{k})$ where $s_{k}\in\{+1,-1\}$, $\delta_{+}\equiv 0$ and $\delta_{-}\equiv 1$. The proof of Eqs. (31) and (32) is given in Appendix H. The commutation relations of $\chi_{\pi}$ with the transfer matrices are evident from Eqs. (31) and (32): $\chi_{\pi}$ commutes with $\tau_{+}(u)$ for odd $N$ but anticommutes with $\tau_{-}(u)$ for even $N$.
@@ V Conclusion
We have established the complete integrability of the clean SYK models with $p$-body interactions by revealing their connection to the critical transverse field Ising chain. The key discovery is that the generalized SYK Hamiltonians emerge from transfer matrices constructed using the R-matrix of the critical Ising chain, which satisfies the Yang-Baxter equation. Our approach provides explicit solutions for all clean SYK models, extending the previously known results [15] for $H_{4}$ and $(H_{3})^{2}$ to the general $p$-body case $H_{p}$. This unexpected connection between the family of clean SYK models and the critical Ising chain places the former within the well-established framework of Yang-Baxter integrability.
Furthermore, our exact solution naturally extends to hybrid models that combine SYK charges with local conserved charges of the critical Ising chain. Such models, with Hamiltonians that are linear combinations of long-range and short-range interaction terms, have also been investigated in the literature [58], where the SYK model and the Kitaev chain are coupled. Our results may provide new insights into the nature of such models.